OpenStudy (anonymous):

this wants to know if there is a solution or not. 4-5 absolute value (-9x-4)=19

OpenStudy (jdoe0001):

4-5| -9x-4|=19 ?

sammixboo (sammixboo):

\(|4-5|(-9\times -4)~=~19\) Tell me what \((-9\times -4)\) equals first :)

OpenStudy (anonymous):

yes

sammixboo (sammixboo):

Are you there?

OpenStudy (anonymous):

postive 36

OpenStudy (anonymous):

can you show me how to do it?

OpenStudy (anonymous):

btw its 4-5|-9x-4|=19